# Enhancing Group Fairness in Federated Learning through Personalization

### v1

@@ Abstract
Personalized Federated Learning (FL) algorithms collaboratively train customized models for each client, enhancing the accuracy of the learned models on the client’s local data (e.g., by clustering similar clients, or by fine-tuning models locally). In this paper, we investigate the impact of such personalization techniques on the group fairness of the learned models, and show that personalization can also lead to improved (local) fairness as an unintended benefit. We begin by illustrating these benefits of personalization through numerical experiments comparing two classes of personalized FL algorithms (clustering and fine-tuning) against a baseline FedAvg algorithm, elaborating on the reasons behind improved fairness using personalized FL, and then providing analytical support. Motivated by these, we further propose a new, Fairness-aware Federated Clustering Algorithm, Fair-FCA, in which clients can be clustered to obtain a (tuneable) fairness-accuracy tradeoff. Through numerical experiments, we demonstrate the ability of Fair-FCA to strike a balance between accuracy and fairness at the client level.
@@ 1 Introduction
Federated Learning (FL) has emerged as a pivotal paradigm for collaboratively training a shared machine learning model across distributed datasets/clients in a privacy-preserving manner (Kairouz et al., 2021). Shared, global models learned through FL can (potentially) outperform standalone models (those trained individually by each client in the absence of collaboration), owing to their ability to effectively aggregate knowledge from diverse clients. Popular FL algorithms such as FedAvg (McMahan et al., 2017) have been found to perform well, particularly when clients have independent and identically distributed (iid) datasets. However, real-world applications often involve heterogeneous datasets across clients, a situation where the performance of FedAvg can considerably deteriorate (Li et al., 2019b). Moreover, the standard FL approach yields a single global model, lacking customization for each client. Consequently, clients with heterogeneous datasets may experience low local accuracy (Li et al., 2020; Tan et al., 2022; Karimireddy et al., 2020). To address these limitations, a spectrum of personalized FL techniques has been proposed, to enhance the local accuracy of the learned models, while keeping some of the benefits of collaborative learning.
@@
In addition to handling issues of data heterogeneity, training FL models that can uphold societal values, as formalized through notions of algorithmic (group) fairness (Barocas et al., 2019), is an important consideration. Consider as an example the scenario where FL is used to train a foundation model/LLM on local datasets, contributed by diverse participants from different regions and communities (Kuang et al., 2023). Without careful oversight, the resulting model will favor language and content preferred by the majority contributors, often disregarding the unique linguistic nuances and cultural contexts of minority groups (Durmus et al., 2023). As another example, existing research (Kirdemir et al., 2021) finds structural and systemic bias in YouTube video recommendation systems. In this case, the collaboratively trained recommendation model begins to favor certain perspectives over others, inadvertently reinforcing biases present in the data. As a result, users are fed information that perpetuates stereotypes, causing harm to underrepresented communities. These examples motivate the development of FL algorithms that can meet certain fairness criteria.
@@
In this paper, we establish an alignment between these two considerations in the developing FL algorithms: personalization can also improve fairness. That is, we show that personalization techniques (which are designed to improve the local accuracy of the learned models) can also enhance algorithmic fairness (i.e., they can reduce the disparities in how the resulting model treats individuals from different protected groups, once the model is deployed by the client).
@@
We begin by illustrating this alignment between personalization and fairness through numerical experiments comparing the average local fairness achieved by two personalized FL algorithms, Clustered-FL (Ghosh et al., 2020) and MAML-FL (Fallah et al., 2020), against FedAvg and standalone learning (where each client learns by itself, and there is no collaboration), under three notions of fairness (Statistical Parity, Equality of Opportunity and Equalized Odds), and using both synthetic and real-world data (different prediction tasks in the “Retiring Adult” dataset (Ding et al., 2021)). Based on these experiments, we highlight two potential factors driving the alignment:
@@
      2. Computational advantages due to alignments in local accuracy and fairness. Moreover, we identify cases where the sample distribution within clients is such that improving accuracy also promotes fairness. More intuitively, consider the concept of clustered FL, a personalized FL technique, where clients are organized into groups based on similarities in model performance (a proxy for data similarities). We argue that this clustering can be seen as a means to foster a model that is both more precise and more equitable, as it effectively treats information originating from other clusters as noise, which, if left unaddressed, would have led to model divergence, potentially compromising both cluster-specific performance and fairness. Building on this intuition, we also provide analytical support of the alignment between personalization and fairness (Propositions 1 and 2).
To the best of our knowledge, this work is the first to identify the (unintended) fairness benefits of personalization techniques in FL. Most existing works primarily focus on refining algorithms to attain either improved local accuracy (the personalized FL literature), or, enhanced (global) fairness (the fair FL literature); we discuss related work from each direction in Section 2. In contrast, our work examines the influence of existing personalization techniques on (local) fairness, and points to inherent features of collaborative learning and personalization that can advance fairness.
Additionally, prior work points to the challenge of balancing the trade-off between fairness and accuracy of an algorithm, as enhancing fairness often comes at the cost of accuracy degradation (Gu et al., 2022; Ezzeldin et al., 2023; Chien & Danks, 2023). In contrast, we identify instances that FL in general, and personalized FL in particular, can improve both accuracy and fairness compared to standalone learning. Furthermore, inspired by our findings, we propose a new, Fairness-aware Federated Clustering Algorithm, Fair-FCA. Our algorithm is an extension of the Clustered-FL (IFCA) algorithm of Ghosh et al. (2020), by modifying the cluster identity assignment step to encompass a fairness criterion. Given our finding that personalization can offer dual benefits, we find that this modification can be used to strike a better balance between accuracy and fairness at the client level.
@@
We conduct extensive numerical experiments, under different notions of fairness, and using both real-world and synthetic data, to show that personalization in FL (which is intended to improve local accuracy) can also improve local fairness as an unintended benefit (Section 4). Accordingly, we highlight the potential statistical and computational reasons leading to this alignment.
2. We analytically show that, under certain conditions, personalization via clustering can lead to improved local fairness compared to a (non-personalized) FedAvg model (Propositions 1 and 2).
@@
3. We propose a Fairness-aware Federated Clustering Algorithm, Fair-FCA, an extension of the IFCA federated clustering algorithm (Ghosh et al., 2020), which takes fairness into account when (iteratively) determining clients’ cluster memberships (Algorithm 1), and show that it can provide desired fairness-accuracy trade-offs by adjusting a hyperparameter (Section 6).
2 Related works
@@
Our work is at the intersection of two literatures: personalized FL, and achieving fairness in FL. Here, we discuss the most closely related works; we review additional related work in Appendix B.
@@
Personalized FL. Existing literature can be categorized based on how personalization is achieved. In this paper, we investigate personalization through clustering and local fine-tuning.
@@
Clustering: Mansour et al. (2020) use a hypothesis-based clustering approach by minimizing the sum of loss over all clusters. Sattler et al. (2020) use the idea that cosine similarity between weight updates of different clients is highly indicative of the similarity of data distribution. Nardi et al. (2022) use a decentralized learning idea by exchanging local models with other clients to find the neighbor/group which has a high accuracy even using other clients’ models. Zheng et al. (2022) learn a weighted and directed graph that indicates the relevance between clients. Ghosh et al. (2020) use a distributed learning idea by broadcasting all clients models to others, and collecting back the cluster identity from clients who can identify good performance when using others’ models.
@@
Local fine-tuning: Fallah et al. (2020) propose using a Model Agnostic Meta Learning (MAML) framework, where clients run additional local gradient steps to personalize the global model. Arivazhagan et al. (2019); Jiang & Lin (2022) propose using deep learning models with a combination of feature extraction layers (base) and global/local head (personalization). Jiang & Lin (2022), inspired by Arivazhagan et al. (2019), further consider robustifying against distribution shifts.
@@
In this paper, we will consider the Ghosh et al. (2020) Clustered-FL (IFCA) algorithm from the clustering category, and the Fallah et al. (2020) MAML-FL algorithm from the local fine-tuning category. Further, our proposed algorithm, Fair-FCA, belongs to the clustering category.
@@
Fairness in FL. This literature, surveyed recently in Shi et al. (2023); Rafi et al. (2023), can also be categorized depending on the adopted notion of fairness. Notably, several works (e.g., (Li et al., 2019a, 2021; Zhang et al., 2021; Wang et al., 2021)) on fair FL consider a notion of performance fairness, assessing whether the learned model(s) achieve uniform accuracy across all clients. In contrast, we focus on notions of social (group) fairness (also studied in, e.g., (Abay et al., 2020; Zhang et al., 2020; Du et al., 2021; Gálvez et al., 2021; Zeng et al., 2021; Ezzeldin et al., 2023)), which compare statistics such as selection rate, true positive rate, or false positive rate, across different demographic/protected groups, to identify any bias in how humans will be treated once the learned models are deployed. Specifically, Abay et al. (2020) propose pre-processing (reweighting with differential privacy) and in-processing (adding fairness-aware regularizer) methods to mitigate biases while protecting the data privacy. Zhang et al. (2020) propose a reinforcement-learning based, fair FL framework, in which the reward is aligned with maximizing the global accuracy while minimizing the discrimination index between groups. Du et al. (2021) propose a fairness-aware FL framework to train a globally fair model through a kernel reweighting technique on both loss function and fairness constraints. Gálvez et al. (2021) introduce an algorithm to enforce group fairness in FL by extending the modified method of differential multiplier to empirical risk minimization with fairness constraints. Ezzeldin et al. (2023) propose fairness-aware aggregation in FL by adjusting the model aggregation weights to prioritize clients whose local fairness is lower than the global fairness. Similarly, Zeng et al. (2021) update model aggregation weights, but by iteratively solving a bi-level optimization problem similar to the centralized fair learning algorithm.
In contrast to all these works, we do not aim to impose a fairness constraint, but show that improved (local) fairness (and a better fairness-accuracy tradeoff) can be achieved either by personalization alone, or by grouping clients into clusters based on a fairness-aware cluster assignment metric.
@@ 3 Problem setting and preliminaries
We study a federated learning setting consisting of $n$ clients. The clients can (potentially) be categorized into two clusters, denoted $c\in\{\mathcal{C}_{\alpha},\mathcal{C}_{\beta}\}$, based on similarities in their data distributions. (We consider $K\geq 2$ clusters in Section 6.) Let $p$ denote the fraction of clients belonging to cluster $\mathcal{C}_{\alpha}$.
@@
Each client $i$ is tasked with a binary classification problem. The client’s dataset consists of samples $z=(x,y,g)$, where $x\in\mathbb{R}$ is a feature, $y\in\{0,1\}$ is the true label of the sample, and $g\in\{a,b\}$ is a demographic group membership based on a protected attribute (e.g., race, gender, or age). Client $i$’s samples are drawn independently at random from joint feature-label-group distributions with pdfs $\{f^{y,c}_{g}(x),y\in\{0,1\},g\in\{a,b\}\}$; note that these distributions are cluster-dependent.
Standalone learning. Clients can use their local data to (independently) learn a threshold-based, binary classifiers $h_{\theta}(x):\mathbb{R}\rightarrow\{0,1\}$, [^1] Our analysis will assume one-dimensional features and threshold classifiers. The former can be viewed as the one-dimensional representation of multi-dimensional features obtained from the last layer outputs of a neural network. For the latter, existing works (Corbett-Davies et al., 2017; Raab & Liu, 2021) show that threshold classifiers are optimal when multi-dimensional features can be properly mapped into a one-dimensional space. under which samples with features $x\geq\theta$ are classified as label 1 (i.e., $\hat{y}(\theta)=1$). Clients choose these thresholds to minimize classification errors. Formally, consider a client $i$ from cluster $c$; let $r^{c}_{g}$ be the fraction of its samples that are from group $g$, and $\alpha^{y,c}_{g}$ be the fraction of its samples that are from group $g$ and have true label $y$. The client chooses its decision threshold ${\theta_{i}^{*}}$ to (empirically) solve the following optimization problem:
$\theta_{i}^{*}=\arg\min_{{\theta}}~{}~{}\sum_{g\in\{a,b\}}r^{c}_{g}\Big{(}\alpha^{1,c}_{g}\int_{-\infty}^{\theta}f^{1,c}_{g}(x)\mathrm{d}x+\alpha^{0,c}_{g}\int_{\theta}^{+\infty}f^{0,c}_{g}(x)\mathrm{d}x\Big{)}~{}.$ (1)
@@
Federated learning. While clients can learn standalone models, Federated learning paradigms enable them to leverage the power of collaborative learning. This is exemplified by the FedAvg algorithm (McMahan et al., 2017), in which the objective is to learn a single global model $\theta$ that minimizes the average loss across all clients. Learning happens over rounds. During each communication round $t$, each client $i$ receives the latest global model $\theta^{t-1}$ from the server, performs a local update on this model based on its local data, and sends back its updated model $\theta^{t}_{i}$ to the server. The server then aggregates all the received local models (through a weighted sum) to create the new global model for the next round. While FedAvg can successfully learn a global model, if the clients are heterogeneous, the local loss of the shared global model may be high locally for each client. This has motivated the development of personalized FL algorithms, two types of which we consider in this work: clustering and local fine-tuning. The Clustered-FL algorithm of Ghosh et al. (2020) alternates between two steps: clustering/grouping clients based on model performance, and optimizing model parameters via gradient descent in each cluster separately, leading to distinct, cluster-specific models. Similarly, the MAML-FL algorithm of Fallah et al. (2020) takes an extra step beyond the FedAvg approach. In it, clients fine-tune the received global model by running additional gradient steps on it locally, to specialize it to their local data and reduce their local loss.
@@
Assessing Fairness. Given the sensitive demographic information $g$ (e.g., race, gender, age) contained in each data point in our problem setting, we further assess the group fairness of the learned models. Different notions of group fairness have been proposed, each assessing a specific statistical disparity in how the model treats individuals from different groups. Here, we consider three notions of fairness: (1) Statistical Parity (SP) (Dwork et al., 2012; Ezzeldin et al., 2023), which compares the difference in selection rate between the two groups (i.e., $\Delta_{\texttt{SP}}(\theta):=|{Pr}(\hat{y}(\theta)=1|g=a)-{Pr}(\hat{y}(\theta)=1|g=b)|$); (2) Equality of Opportunity (EqOp) (Hardt et al., 2016), which compares the differences in true positive rates across groups (i.e., $\Delta_{\texttt{EqOp}}(\theta):=|Pr(\hat{y}(\theta)=1|g=a,y=1)-{Pr}(\hat{y}(\theta)=1|g=b,y=1)|)$; and (3) Equalized Odd (EO) (Hardt et al., 2016), which is set to either the true positive rate or false positive rate difference between the two protected groups, whichever larger (i.e., $\Delta_{\texttt{EO}}(\theta):=\max_{i\in\{0,1\}}|{Pr}(\hat{y}(\theta)=1|g=a,y=i)-{Pr}(\hat{y}(\theta)=1|g=b,y=i)|$).
@@
Comparing different algorithms. To compare the different algorithms’ performance, we will contrast the (average of the) local accuracy of each client’s local model (which can be its standalone model, the global FedAvg model, a cluster-specific model, or a fine-tuned model) when applied to the client’s local data. We will similarly contrast the different algorithms in terms of the (average of) local $f$-fairness, evaluating each client’s local model on its local data under $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$.
4 Personalization can also improve fairness: numerical illustration

### v2

@@ Abstract
Personalized Federated Learning (FL) algorithms collaboratively train customized models for each client, enhancing the accuracy of the learned models on the client’s local data (e.g., by clustering similar clients, by fine-tuning models locally, or by imposing regularization terms). In this paper, we investigate the impact of such personalization techniques on the group fairness of the learned models, and show that personalization can also lead to improved (local) fairness as an unintended benefit. We begin by illustrating these benefits of personalization through numerical experiments comparing several classes of personalized FL algorithms against a baseline FedAvg algorithm, elaborating on the reasons behind improved fairness using personalized FL, and then providing analytical support. Motivated by these, we then show how to build on this (unintended) fairness benefit, by further integrating a fairness metric into the cluster-selection procedure of clustering-based personalized FL algorithms, and improve the fairness-accuracy trade-off attainable through them. Specifically, we propose two new fairness-aware federated clustering algorithms, Fair-FCA and Fair-FL+HC, extending the existing IFCA and FL+HC algorithms, and demonstrate their ability to strike a (tuneable) balance between accuracy and fairness at the client level.
@@ 1 Introduction
Federated Learning (FL) has emerged as a pivotal paradigm for collaboratively training a shared machine learning model across distributed datasets/clients in a privacy-preserving manner (kairouz2021advances). Shared, global models learned through FL can (potentially) outperform standalone models (those trained individually by each client in the absence of collaboration), owing to their ability to effectively aggregate knowledge from diverse clients. Popular FL algorithms such as FedAvg (mcmahan2017communication) have been found to perform well, particularly when clients have independent and identically distributed (iid) datasets. However, real-world applications often involve heterogeneous datasets across clients, a situation where the performance of FedAvg can considerably deteriorate (li2019convergence). Moreover, the standard FL approach yields a single global model, lacking customization for each client. Consequently, clients with heterogeneous datasets may experience low local accuracy (li2020federated; tan2022towards; karimireddy2020scaffold). To address these limitations, a spectrum of personalized FL techniques has been proposed, to enhance the local accuracy of the learned models, while keeping some of the benefits of collaborative learning.
@@
Aside from handling issues of data heterogeneity, training FL models that can uphold societal values, as formalized through notions of algorithmic (group) fairness (barocas-hardt-narayanan), has been a focus of recent research. Consider as an example the scenario where FL is used to train a foundation model/LLM on local datasets, contributed by diverse participants from different regions and communities (kuang2023federatedscope). Without careful oversight, the resulting model will favor language and content preferred by the majority contributors, often disregarding the unique linguistic nuances and cultural contexts of minority groups (durmus2023towards). As another example, existing research (kirdemir2021assessing) finds structural and systemic bias in YouTube video recommendation systems. In this case, the collaboratively trained recommendation model begins to favor certain perspectives over others, inadvertently reinforcing biases present in the data. As a result, users are fed information that perpetuates stereotypes, causing harm to underrepresented communities. These examples motivate the development of FL algorithms that can meet certain fairness criteria.
@@
In this paper, we establish an alignment between these two considerations in developing FL algorithms: personalization can also improve fairness. That is, we show that personalization techniques (which are designed to improve the local accuracy of the learned models) can also enhance algorithmic fairness as an unintended benefit (i.e., they can reduce the disparities in how the resulting model treats individuals from different protected groups, once the model is deployed by the client).
@@
We begin by illustrating this alignment between personalization and fairness through numerical experiments comparing the average local fairness achieved by several personalized FL algorithms (IFCA (ghosh2020efficient), MAML-FL (fallah2020personalized), FedProx (li2020federated), FL+HC (briggs2020federated), and CFL-Cosine (sattler2020clustered)), against FedAvg (mcmahan2017communication) and standalone learning (where each client learns by itself, and there is no collaboration), under three notions of fairness (Statistical Parity, Equality of Opportunity and Equalized Odds), and using both synthetic and real-world data (“Adult” dataset (Dua:2019), and combinations of two different prediction tasks and two protected attributes in the “Retiring Adult” dataset (ding2021retiring)). Based on these experiments, we highlight two potential factors driving the alignment:
@@
      2. Computational advantages due to alignments in local accuracy and fairness. Moreover, we identify cases where the sample distribution within clients is such that improving accuracy also promotes fairness. More intuitively, consider the concept of clustered FL, a personalized FL technique, where clients are organized into groups based on similarities in model performance (a proxy for data similarities). We argue that this clustering can be seen as a means to foster a model that is both more precise and more equitable, as it effectively treats information originating from other clusters as noise, which, if left unaddressed, would have led to model divergence. (Building on this intuition, we also provide analytical support under certain conditions in Propositions LABEL:prop3 and LABEL:prop1 in Appendix LABEL:sec:analysis).
To the best of our knowledge, this work is the first to identify the (unintended) fairness benefits of personalization in FL. Most existing works primarily focus on refining algorithms to attain either improved local accuracy (the personalized FL literature), or, enhanced (global) fairness (the fair FL literature); we discuss related work from each direction in Section 2. In contrast, our work examines the influence of existing personalization techniques on (local) fairness, and points to inherent features of collaborative learning and personalization that can advance fairness. Additionally, prior work points to the challenge of balancing the trade-off between fairness and accuracy of an algorithm, as enhancing fairness often comes at the cost of lower accuracy (gu2022privacy; ezzeldin2023fairfed; chien2023fairness). In contrast, we identify instances that FL in general, and personalized FL in particular, can improve both accuracy and fairness compared to standalone learning.
Furthermore, inspired by our findings, we propose two new fairness-aware federated clustering algorithms, Fair-FCA and Fair-FL+HC, based on the existing IFCA (ghosh2020efficient) and FL+HC (briggs2020federated) personalized FL algorithms, respectively. Both Fair-FCA and Fair-FL+HC algorithms modify the cluster identity assignment step to encompass a fairness criterion. Building on our finding that personalization can offer dual benefits, we demonstrate that this additional modification can be used to strike a better balance between accuracy and fairness at the client level.
@@
We conduct extensive numerical experiments, under different notions of fairness, and using both real-world and synthetic data, to show that personalization in FL (which is intended to improve local accuracy) can also improve local fairness as an unintended benefit (Section 4). We highlight the potential statistical and computational reasons leading to this alignment, and provide analytical support under certain conditions (Prop. LABEL:prop3 and LABEL:prop1).
@@
2. New fairness-aware, personalized federated clustering algorithms. We propose two new algorithms, Fair-FCA and Fair-FL+HC, by modifying the existing IFCA (ghosh2020efficient) and FL+HC (briggs2020federated) personalized FL algorithms, to take fairness into account when (iteratively) determining clients’ cluster memberships (Algorithms 1 and 2). We show that this can provide desired fairness-accuracy trade-offs by adjusting a hyperparameter (Section 5), and elaborate on when/how this technique can be applied to other (clustering-based) personalized FL methods.
2 Related work
@@
Our work is at the intersection of two literatures: personalized FL, and achieving fairness in FL. Here, we discuss the most closely related works; we review additional related work in Appendix LABEL:app:related.
@@
Personalized FL. Existing literature can be categorized based on how personalization is achieved. Here, we investigate personalization through clustering, local fine-tuning, and regularization.
@@
Clustering: mansour2020three use a hypothesis-based clustering approach by minimizing the sum of loss over all clusters. sattler2020clustered use the idea that cosine similarity between weight updates of different clients is highly indicative of the similarity of data distribution. nardi2022anomaly use a decentralized learning idea by exchanging local models with other clients to find the neighbor/group which has a high accuracy even using other clients’ models. zheng2022diple learn a weighted and directed graph that indicates the relevance between clients. ghosh2020efficient use a distributed learning idea by broadcasting all clients models to others, and collecting back the cluster identity from clients who can identify good performance when using others’ models.
@@
Local fine-tuning: fallah2020personalized propose using a Model Agnostic Meta Learning (MAML) framework, where clients run additional local gradient steps to personalize the global model. arivazhagan2019federated; jiang2022test propose using deep learning models with a combination of feature extraction layers (base) and global/local head (personalization). jiang2022test, inspired by arivazhagan2019federated, further consider robustifying against distribution shifts.
Model regularization: hanzely2020federated; sahu2018convergence; li2020federated; li2021ditto add a regularization term with mixing parameters to penalize the distance between the local and global models. In particular, sahu2018convergence has a pre-set regularization parameter and allows for system heterogeneity. li2021ditto consider improving accuracy, while being robust to data and model poisoning attacks, and fair. Similarly, t2020personalized formulate a bi-level optimization problem, which helps decouple personalized model optimization from learning the global model. huang2021personalized propose the FedAMP algorithm which also introduces an additional regularization term, but differs from previous works in that they encourage similar clients to collaborate more.
@@
In this paper, we will consider three variants of clustered FL algorithms (IFCA (ghosh2020efficient), FL+HC (briggs2020federated), and CFL-Cosine (sattler2020clustered)) from the clustering category. Additionally, we consider MAML-FL by fallah2020personalized from the local fine-tuning category, and FedProx by li2020federated from the model regularization category. Our new proposed algorithms, Fair-FCA and Fair-FL+HC, will belong to the clustering category.
@@
Fairness in FL. This literature, surveyed recently in shi2023towards; rafi2023fairness, can also be categorized depending on the adopted notion of fairness. Notably, several works (e.g., (li2019fair; li2021ditto; zhang2021unified; wang2021federated)) on fair FL consider a notion of performance fairness, assessing whether the learned model(s) achieve uniform accuracy across all clients. In contrast, we focus on notions of social (group) fairness (also studied in, e.g., (abay2020mitigating; zhang2020fairfl; du2021fairness; galvez2021enforcing; zeng2021improving; ezzeldin2023fairfed)), which compare statistics such as selection rate, true positive rate, or false positive rate, across different demographic/protected groups, to identify any bias in how humans will be treated once the learned models are deployed. In contrast to all these works, we do not aim to impose a fairness constraint during training or aggregation of the (global) models, but show that improved (local) fairness (and a better fairness-accuracy tradeoff) can be achieved either by personalization alone, or by grouping clients into clusters based on a fairness-aware cluster assignment metric.
@@ 3 Problem setting and preliminaries
We study a federated learning setting consisting of $n$ clients. Each client $i$ is tasked with a binary classification problem. The client’s dataset consists of samples $z=(x,y,g)$, where $x\in\mathbb{R}^{n}$ is a feature, $y\in\{0,1\}$ is the true label of the sample, and $g\in\{a,b\}$ is a demographic group membership based on a protected attribute (e.g., race, gender, or age). Client $i$’s samples are drawn independently at random from joint feature-label-group distributions with pdfs $\{f^{y,i}_{g}(x),y\in\{0,1\},g\in\{a,b\}\}$; note that these distributions are client-dependent and potentially different, motivating the need for personalization of the learned models. Let $\mathcal{D}_{i}$ denote the local dataset of client $i$.
Standalone learning. Clients can use their local data $\mathcal{D}_{i}$ to (independently) learn binary classifiers $h_{\theta_{i}}(x):\mathbb{R}^{n}\rightarrow\{0,1\}$, where $\theta_{i}$ are the parameters of the classifier (e.g., weights of a NN).
@@
Federated learning. While clients can learn standalone models, Federated Learning (FL) paradigms enable them to leverage the power of collaborative learning. This is exemplified by the FedAvg algorithm (mcmahan2017communication), in which the objective is to learn a single global model $\theta$ that minimizes the average loss across all clients. Learning happens over rounds. During each communication round $t$, each client $i$ receives the latest global model $\theta^{t-1}$ from the server, performs a local update on this model based on its local data $\mathcal{D}_{i}$, and sends back its updated model $\theta^{t}_{i}$ to the server. The server then aggregates all the received local models (through a weighted sum) to create the new global model for the next round. While FedAvg can successfully learn a global model, if the clients’ datasets $\mathcal{D}_{i}$ are heterogeneous, the local loss of the shared global model may be high locally for each client. This has motivated the development of personalized FL algorithms, five of which we consider in this work: clustering-based methods (specifically, IFCA, FL+HC, and CFL-Cosine), a local fine-tuning method (MAML-FL), and one that imposes model regularization (FedProx).
Personalized Federated learning. We provide a high-level description of the personalized FL methods considered in this paper. The IFCA algorithm of ghosh2020efficient alternates between two steps: clustering/grouping clients based on model performance, and optimizing model parameters via gradient descent in each cluster separately, leading to distinct, cluster-specific models. The FL+HC algorithm of briggs2020federated uses a hierarchical clustering technique to group clients while minimizing intra-cluster variance, measured by the Euclidean distance of their models. The CFL-Cosine algorithm of sattler2020clustered bipartitions clients, aiming to minimize the maximum cosine similarity of clients’ gradient updates. The MAML-FL algorithm of fallah2020personalized builds upon the FedAvg approach by allowing clients to further fine-tune the received global model through additional local gradient steps, thereby specializing it to their data and reducing local loss. Similarly, the FedProx algorithm of li2020federated introduces a $l_{2}$ regularization term for each client to learn a mix of local and global models.
@@
Assessing Fairness. Given the sensitive demographic information $g$ (e.g., race, gender, age) contained in each data point in our problem setting, we further assess the group fairness of the learned models. Different notions of group fairness have been proposed, each assessing a specific statistical disparity in how the model treats individuals from different groups. Here, we consider three notions of fairness: (1) Statistical Parity (SP) (dwork2012fairness; ezzeldin2023fairfed), which compares the difference in selection rate between the two groups (i.e., $\Delta_{\texttt{SP}}(\theta):=|{Pr}(\hat{y}(\theta)=1|g=a)-{Pr}(\hat{y}(\theta)=1|g=b)|$, where $\hat{y}(\theta)$ is the label assigned to data points under model $\theta$); (2) Equality of Opportunity (EqOp) (hardt2016equality), which compares the differences in true positive rates across groups (i.e., $\Delta_{\texttt{EqOp}}(\theta):=|Pr(\hat{y}(\theta)=1|g=a,y=1)-{Pr}(\hat{y}(\theta)=1|g=b,y=1)|)$; and (3) Equalized Odd (EO) (hardt2016equality), which is set to either the true positive rate or false positive rate difference between the two protected groups, whichever larger (i.e., $\Delta_{\texttt{EO}}(\theta):=\max_{i\in\{0,1\}}|{Pr}(\hat{y}(\theta)=1|g=a,y=i)-{Pr}(\hat{y}(\theta)=1|g=b,y=i)|$).
@@
Comparing different algorithms. To compare the different algorithms’ performance, we will contrast the (average of the) local accuracy of each client’s local model $\theta_{i}$ (which can be obtained by finding its standalone model, the global FedAvg model, a cluster-specific local model, or a fine-tuned local model, or a regularized local model) when applied to the client’s local data $\mathcal{D}_{i}$. We will similarly contrast the different algorithms in terms of the (average of) local $f$-fairness, evaluating each client’s local model on its local data under $f\in\{\texttt{SP},\texttt{EqOp},\texttt{EO}\}$.
4 Personalization can also improve fairness